# Current-induced molecular dissociation: Topological insulators as robust reaction platforms

### v1

@@ I Introduction
Figure 1: (a) Scheme of the catalysis model, based on the interaction between the adsorbate (red circles) and the adsorbent, composed of an armchair graphene nanoribbon. The contacts extend on both side of the sample and are represented by yellow regions. The shaded region around the adsorbate indicates the extent of the coupling with the adsorbent. (b) Honeycomb lattice with armchair edge termination. The shaded region denotes the unit cell, containing $2N$ carbon atoms.
@@
Since transport phenomena represent one of the most promising features of topological insulators, in this work we investigate the role of topological states in the destabilization of molecules interacting with edge currents within a transport setup. Modeling a full 3D catalytic system in non-equilibrium conditions at the quantum transport level is computationally challenging; therefore, to capture the essential physics while retaining numerical tractability, we restrict our analysis to two-dimensional substrates. In this framework, the catalytic device is modeled as a nanoribbon connected to semi-infinite leads, with the molecule placed near one of its edges to mimic surface adsorption in a real 3D catalyst [see Fig. 1(a)]. In addition, we use a single-particle approach based on tight-binding (TB) models for the composite system. Although neglecting many-body interactions would lead to some quantitative limitation, the obtained results in this work illustrate the role of TIs, which could be extended to a more involved description of the physical system. This minimal setup thus allows us to explore current-induced effects on molecular dissociation while retaining the key differences between standard metallic and topological substrates.
@@
As a proof of concept, we describe a diatomic molecule as a two-orbital system. Applying a bias voltage across the leads drives a current through the substrate, enabling us to study how charge flow modifies the electronic occupation of molecular levels and the resulting electronic forces. To establish a consistent comparison between metallic and topological catalysts in terms of efficiency and robustness of current-induced molecular destabilization, we consider a graphene ribbon with armchair edge termination. By introducing a spin-orbit coupling (SOC) term, the ribbon becomes a topological insulator described by the Kane-Mele model [21, 22], opening a topological gap that supports robust helical edge states. In this framework, for both substrates we show that a finite bias can destabilize molecular bonds by promoting electron transfer from bonding to antibonding levels, which counteracts the equilibrium attractive force of the molecular bond, thereby enhancing current-induced dissociation. Comparing trivial graphene with topological substrates, we find that extended states in graphene reduce molecular occupancies of the molecular states as the ribbon width increases, while localized edge states in topological insulators preserve them, sustaining catalytic efficiency. Furthermore, we demonstrate that topological substrates display a remarkable robustness against disorder, in contrast to the strong sensitivity observed in graphene. These results highlight the potential of topological edge states as efficient and resilient platforms for current-driven catalysis.
@@ II.2 Theoretical framework: NEGF approach
To compute the studied physical properties, we employ the non-equilibrium Green’s function (NEGF) formalism [29, 30] which, in the approximation of non-interacting electrons, offers computational efficiency compared to directly solving the Schrödinger equation under non-equilibrium conditions. Within this formalism, the non-equilibrium stationary density matrix can be computed as $\bm{\rho}(t)=-i\hbar\bm{G}^{<}(t)$, where $\bm{G}^{<}(t)$ is the lesser Green’s function:
$\bm{G}^{<}(t)=\int\frac{\mathrm{d}\epsilon}{2\pi\hbar}\bm{G}^{r}(\epsilon)\bm{\Sigma}^{<}(\epsilon)\bm{G}^{a}(\epsilon),$ (10)
@@
$\bm{G}^{r}(\epsilon)=\lim_{\eta\rightarrow 0^{+}}\left[\left(\epsilon+i\eta\right)\bm{I}-\bm{H}\right]^{-1}.$ (12)
Given the infinite-dimensional Hilbert space associated with the contacts, a direct calculation of the above Green’s function matrix results unfeasible. Nevertheless, an effective Hamiltonian $\bm{H}_{\mathrm{eff}}$ can be obtained by including the retarded self-energy corrections $\bm{\Sigma}^{r}_{\alpha}$ which account for the presence of the leads. The calculation of the self-energies exploits the translational invariance within the leads, yielding a fast decimation procedure [31, 32].
@@
where we omit the time argument in $\bm{\rho}$ because we are only interested in the stationary solution. In this expression, an “occupation kernel” can be defined as $\bm{\mathcal{K}}^{\rho}_{\alpha}=\bm{G}^{r}\bm{\Gamma}_{\alpha}\bm{G}^{a}/\pi$, which encapsulates the contribution to the density matrix due to electrons incoming from the $\alpha$ lead. It may also be convenient to separate the equilibrium and non-equilibrium contributions in Eq. (13). By rewriting the distribution functions as $f_{\alpha}=f_{0}+(f_{\alpha}-f_{0})$, we obtain $\bm{\rho}=\bm{\rho}^{\mathrm{eq}}+\bm{\rho}^{\mathrm{ne}}$, where the non-equilibrium term is given by:
@@
Regarding the total Hamiltonian, the electron force $F$ can also be decomposed into a contribution $F_{\mathrm{s}}$ arising from the interaction with the substrate through $\hat{H}_{\mathrm{int}}$ and a contribution $F_{\mathrm{d}}$ associated with the molecule itself through $\hat{H}_{\mathrm{d}}$. The former involves the hybridization between the molecular states and the lattice sites, while the latter accounts for the molecular occupancies. It is found that, in general, $F_{\mathrm{s}}\ll F_{\mathrm{d}}$. Throughout the rest of this work, we will restrict our analysis to the intramolecular term $F_{\mathrm{d}}$, defined through the following force operator:
@@
To compute the electronic occupation of the molecular levels, we project the full system onto the molecular subspace. In particular, we evaluate the occupancies in a renormalized eigenbasis that accounts for the presence of the substrate. To this end, we construct an effective Hamiltonian for the molecular subspace that includes the renormalization induced by the substrate degrees of freedom:
$\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)=\bm{H}_{\mathrm{dd}}+\bm{V}_{\mathrm{ds}}\left(\epsilon\bm{I}-\bm{H}_{\mathrm{ss}}\right)^{-1}\bm{V}_{\mathrm{sd}},$ (17)
where $\bm{H}_{\mathrm{dd}}$ and $\bm{H}_{\mathrm{ss}}$ are the bare molecular and substrate Hamiltonians, respectively, while $\bm{V}_{\mathrm{ds}}$ and $\bm{V}_{\mathrm{sd}}$ are the coupling matrices between these subsystems. From the transformation matrix $\bm{U}(\epsilon)$ that diagonalizes $\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)$, [^1] We diagonalize the effective Hamiltonian by taking $\epsilon$ as a fixed parameter, according to $\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)\bm{u}(\epsilon)=\lambda(\epsilon)\bm{u}(\epsilon)$, i.e., we do not solve the stationary Schrödinger equation which leads to the poles of the Green’s function $\tilde{\bm{G}}^{r}_{\mathrm{dd}}(\epsilon)=[(\epsilon+i0^{+})\bm{I}_{\mathrm{dd}}-\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)]^{-1}$ in the molecular subspace. we compute the density matrix in this reduced basis through the kernel:
@@
where $\bm{\Pi}_{\mathrm{d}}$ is the projector operator in the molecular subspace. This transformation provides direct access to the occupancies of the substrate-renormalized molecular states. We will call these ‘bonding’ (b) and ‘antibonding’ (a) states, since they essentially preserve the properties of the symmetric and antisymmetric states defined in Eq. (6), for the considered regimes of parameters.
@@ III.1 Non-equilibrium molecular destabilization
We initiate our analysis by considering an armchair-edged graphene nanoribbon defined by $N$ carbon atoms along the transversal direction [36] [see Fig. 1(b)]. In particular, we set $N$ under the condition $N=3\ell+2$, with $\ell\in\mathbb{N}$, in order to obtain a metallic ribbon. For this analysis we use $\ell=25$, resulting in a nanoribbon width of approximately $W=\sqrt{3}(N-1)a/2\approx 9.3$ nm. The molecular center is located at a distance of $1.5$ Å above the substrate plane, with its bond axis aligned along $\theta=\pi/2$, i.e., parallel to the $y$ direction, as illustrated in Fig. 1(a). Additionally, its center is located at the lower edge, i.e., $y_{0}=(1-N)\sqrt{3}a/4$, measured from the center of the ribbon. The two atoms are separated by a distance of $1.5$ Å from each other along the bond axis, such that the radius of the molecule is $R=0.75$ Å. For the topological case ($\gamma_{1}\neq 0$), this setup allows us to probe the interaction between the molecule and the localized edge-states. To ensure that the edge states are well defined and spatially separated, the ribbon width must be chosen to be larger than the localization length to avoid hybridization between opposite edges.
Figure 2: Molecular destabilization under non-equilibrium conditions. (a) Occupation probabilities of the bonding (solid lines) and antibonding (dotted lines) levels as functions of the bias voltage for graphene (G, blue lines) and Kane-Mele model (KM, red lines) substrates. (b) Corresponding non-equilibrium intramolecular force, according to Eq. (16). We use a relaxation energy $\eta=10^{-5}\gamma_{0}$, an equilibrium chemical potential $\mu_{0}=0$, and $k_{\mathrm{B}}T=10^{-3}\,\gamma_{0}$.
@@
With these parameters, in our system the bonding level lies below the Dirac point ($\epsilon=0$), and the antibonding level above it. If the molecule were isolated, both levels would be symmetrically placed with respect to the Dirac point. However, due to hybridization with the substrate, this balance is broken and the bonding level shifts more than the antibonding one, ending up closer to the Dirac point. A similar effect could also be achieved by shifting the average molecular energy $(E_{1}+E_{2})/2$ with respect to the Dirac point.
To account for additional relaxation processes due to the environment, we introduce a small imaginary part $\eta$ in the energy variable, i.e., $\epsilon\rightarrow\epsilon+i\eta$. This would correspond, e.g., to coupling the system to a phononic bath that allows electrons to decay into lower energy levels. The value of $\eta$ effectively sets a timescale for the electronic relaxation: a smaller $\eta$ corresponds to a longer lifetime of electronic states, while a larger $\eta$ implies faster relaxation. This allows us to capture the effect of finite-lifetime processes, such as electron-phonon interactions, without the need to model them explicitly (see Appendix A).
@@
In Fig. 2(a) we show the occupation probability of the molecular levels as a function of the applied bias for both trivial graphene ($\gamma_{1}=0$) and topological ($\gamma_{1}\neq 0$) substrates. At equilibrium ($V=0$), the bonding level shows its maximum occupancy while the antibonding level is unoccupied. This is because we set the equilibrium chemical potential at the Dirac point, i.e., $\mu_{0}=0$, and the bonding (antibonding) level lies below (above) this energy point. When a bias voltage is applied and increased, the occupancy of the bonding level starts to decrease while the antibonding level becomes progressively populated. In this case, the molecular levels enter the region known as the “bias window”—that is, between the two chemical potentials. Here, a finite electron current flows through the molecule, thereby modifying the previous equilibrium occupancies.
@@
As can be seen, the variations in the bonding and antibonding occupancies take place at different voltage values: the bonding level depopulates first, while the antibonding occupancy begins to increase at slightly higher voltages. This difference arises from the asymmetric energy alignment of the bonding and antibonding levels with respect to $\mu_{0}$, as previously discussed. Importantly, both the depopulation of the bonding level and the subsequent population of the antibonding level contribute to the destabilization of the molecule. In particular, for large biases the antibonding occupancy can even surpass that of the bonding level. These results indicate that the interaction with either graphene or Kane-Mele substrates can drive molecular dissociation through electron transport. Note that, since the molecule is coupled to an open system, the corresponding occupancies are not necessarily normalized. These deviate from their ideal values because we impose a finite relaxation time through the $\eta$ parameter which, in turn reflects the decay rates to the contacts (see Appendix A). For example, the kernel $\bm{\mathcal{K}}_{\mathrm{L}}^{\rho}]_{\mathrm{b},\mathrm{b}}$, related to the bonding level’s occupation from the left contact, develops a Lorentzian peak shape due to the molecule-substrate hybridization. Its width is characterized by the sum of the rate $\Gamma_{\mathrm{L},\mathrm{b}}$, associated with the molecular hybridization with the left contact, and $\eta$. Energy integration of this kernel [c.f., Eq. (13)] yields a factor $\Gamma_{\mathrm{L},\mathrm{b}}/(\Gamma_{\mathrm{L},\mathrm{b}}+\eta)$ which bounds the level occupancy below its ideal value. Therefore, the $\eta$ parameter establishes an additional decay mechanism that competes with that related with the coupling to the contacts. Indeed, if we set $\eta=0$, the level occupancy recovers its ideal value, regardless of the value of $\Gamma_{\mathrm{L},\mathrm{b}}$.
@@
It is often more intuitive to analyze the destabilization of the dimer in terms of the non-equilibrium force acting on the molecular bond. For this purpose we show, in Fig. 2(b), the non-equilibrium component of the intramolecular force of Eq. (16). In the absence of an applied bias, the non-equilibrium force is zero, which signals molecular stability. As the bias increases, the force exhibits a sharp rise, followed by a small shoulder over a narrow range of $V$ before rising again and reaching its maximum value at higher biases. Once more, this non-monotonic behavior, particularly the presence of the intermediate plateau, can be attributed to the shifts of the molecular levels when coupled with the substrate. Consequently, the initial increase in force is associated with the depletion of the bonding occupancy, while the subsequent rise is driven by the occupation of the antibonding level.
@@
It may be useful to establish a comparison between the equilibrium and non-equilibrium components of the force when the two molecular levels lie within the bias window. By relating the intramolecular force with the occupancies through (see Appendix A):
$F_{\mathrm{d}}=2\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}R}(\rho_{\mathrm{b}}-\rho_{\mathrm{a}}),$ (20)
@@
we obtain, for $eV>2\gamma_{\mathrm{d}}$, that $|F^{\mathrm{ne}}/F^{\mathrm{eq}}|=(1+\rho_{\mathrm{a}}/\rho_{\mathrm{b}})/2$. For the considered substrates we find that $\rho_{\mathrm{a}}>\rho_{\mathrm{b}}$, which implies that the destabilizing non-equilibrium component is larger than the stabilizing equilibrium term. In addition, it can be seen that $F_{\mathrm{d},\mathrm{KM}}^{\mathrm{ne}}>F_{\mathrm{d},\mathrm{G}}^{\mathrm{ne}}$, indicating a higher dissociation efficiency in the topological substrate.
@@
In this section, we analyze the effect of the ribbon width on the molecular occupation, considering the extended nature of metallic states in graphene-based substrates and the localized nature of edge states in topological substrates. This analysis is particularly relevant in systems with coupling to an external environment—modeled via a finite broadening parameter $\eta$—that introduces escape channels that affect the molecular occupancies. A very small $\eta$ would act as a mathematical regularization which, for energy levels below the Fermi energy, yields occupancies close to its ideal value regardless of the system size. A finite $\eta$, on the other hand, allows the ribbon’s spatial characteristics to play a physical role in the transport processes, since the decay rates $\Gamma_{\alpha,\mathrm{b/a}}$ modulate the molecular occupancies. This effect is illustrated in Appendix B.

### v2

@@ I Introduction
Figure 1: (a) Scheme of the catalysis model, based on the interaction between the adsorbate (red circles) and the adsorbent, composed of an armchair graphene nanoribbon with passivated vacancies. The contacts extend on both sides of the sample and are represented by yellow regions. The shaded region around the adsorbate indicates the extent of the coupling with the adsorbent. Inset: Zoom of the molecule-substrate interaction region; with $\gamma_{0}$ the carbon-carbon hopping, $\gamma_{\mathrm{d}}$ the intramolecular coupling, and $\gamma(\ell)$ the hopping term between the atoms of the molecule and the carbon atoms of the substrate. (b) Honeycomb lattice with armchair edge termination. The shaded region denotes the unit cell, containing $2N$ carbon atoms.
@@
Since transport phenomena represent one of the most promising features of topological insulators, in this work we investigate the role of topological states in the destabilization of molecules interacting with edge currents within a transport setup. In this regard, the effect of electronic currents on the potential dissociation of molecules near topological materials is an incipient subject of study. There are, nevertheless, reports in the literature that are related to exchange effects between a nanoparticle (or molecule) and a topological substrate. For example, several works address the role of spin-momentum locking effect in the properties of a nanomagnet weakly coupled to the helical edge states of a 2D TI in a transport setup [21, 22, 23]. In a related direction, in Ref. [24] the authors explore the quantum drag force on a metallic nanoparticle moving above a 2D topological substrate.
@@
For our purposes, modeling a full 3D catalytic system in non-equilibrium conditions at the quantum transport level is computationally challenging; therefore, to capture the essential physics while retaining numerical tractability, we restrict our analysis to two-dimensional substrates. In this framework, the catalytic device is modeled as a nanoribbon connected to semi-infinite leads, with the molecule placed near one of its edges to mimic surface adsorption in a real 3D catalyst [see Fig. 1(a)]. In addition, we use a single-particle approach based on tight-binding (TB) models for the composite system. Although neglecting many-body interactions would lead to some quantitative limitation, the obtained results in this work illustrate the role of TIs, which could be extended to a more involved description of the physical system. This minimal setup thus allows us to explore current-induced effects on molecular dissociation while retaining the key differences between standard metallic and topological substrates.
@@
As a proof of concept, we describe a diatomic molecule as a two-orbital system. Applying a bias voltage across the leads drives a current through the substrate, enabling us to study how charge flow modifies the electronic occupation of molecular levels and the resulting electronic forces. To establish a consistent comparison between metallic and topological catalysts in terms of efficiency and robustness of current-induced molecular destabilization, we consider a graphene ribbon with armchair edge termination. By introducing a spin-orbit coupling (SOC) term, the ribbon becomes a topological insulator described by the Kane-Mele model [25, 26], opening a topological gap that supports robust helical edge states. In this framework, for both substrates we show that a finite bias can destabilize molecular bonds by promoting electron transfer from bonding to antibonding levels, which counteracts the equilibrium attractive force of the molecular bond, thereby enhancing current-induced dissociation. Comparing trivial graphene with topological substrates, we find that extended states in graphene reduce molecular occupancies of the molecular states as the ribbon width increases, while localized edge states in topological insulators preserve them, sustaining catalytic efficiency. Furthermore, we demonstrate that topological substrates display a remarkable robustness against vacancy disorder, in contrast to the strong sensitivity observed in graphene. These results highlight the potential of topological edge states as efficient and resilient platforms for current-driven catalysis.
@@ II.2 Theoretical framework: NEGF approach
To compute the studied physical properties, we employ the non-equilibrium Green’s function (NEGF) formalism [33, 34] which, in the approximation of non-interacting electrons, offers computational efficiency compared to directly solving the Schrödinger equation under non-equilibrium conditions. Within this formalism, the non-equilibrium stationary density matrix can be computed as $\bm{\rho}(t)=-i\hbar\bm{G}^{<}(t)$, where $\bm{G}^{<}(t)$ is the lesser Green’s function. In the stationary limit this function becomes time-independent, and can be obtained as:
$\bm{G}^{<}=\int\frac{\mathrm{d}\epsilon}{2\pi\hbar}\bm{G}^{r}(\epsilon)\bm{\Sigma}^{<}(\epsilon)\bm{G}^{a}(\epsilon),$ (10)
@@
$\bm{G}^{r}(\epsilon)=\lim_{\eta\rightarrow 0^{+}}\left[\left(\epsilon+i\eta\right)\bm{I}-\bm{H}_{\mathrm{sys}}-\bm{\Sigma}^{r}(\epsilon)\right]^{-1},$ (12)
where $\bm{H}_{\mathrm{sys}}$ denotes the central system Hamiltonian including the substrate, the molecule, and the interaction between them. $\bm{\Sigma}^{r}=\bm{\Sigma}^{r}_{\mathrm{L}}+\bm{\Sigma}^{r}_{\mathrm{R}}$ is the retarded self-energy correction due to the contacts, whose calculation exploits the translational invariance within the leads, yielding a fast decimation procedure [35, 36].
@@
where, as in Eq. (10), we omit the time argument because we are only interested in the stationary solution. In this expression, an “occupation kernel” can be defined as $\bm{\mathcal{K}}^{\rho}_{\alpha}=\bm{G}^{r}\bm{\Gamma}_{\alpha}\bm{G}^{a}/\pi$, which encapsulates the contribution to the density matrix due to electrons incoming from the $\alpha$ lead. It may also be convenient to separate the equilibrium and non-equilibrium contributions in Eq. (13). By rewriting the distribution functions as $f_{\alpha}=f_{0}+(f_{\alpha}-f_{0})$, we obtain $\bm{\rho}=\bm{\rho}^{\mathrm{eq}}+\bm{\rho}^{\mathrm{ne}}$, where the non-equilibrium term is given by:
@@
Regarding the total Hamiltonian, the electron force $F$ can also be decomposed into a contribution $F_{\mathrm{s}}$ arising from the interaction with the substrate through $\hat{H}_{\mathrm{int}}$ and a contribution $F_{\mathrm{d}}$ associated with the molecule itself through $\hat{H}_{\mathrm{d}}$. The former involves the hybridization between the molecular states and the lattice sites, while the latter accounts for the molecular occupancies. For the considered parameters, we find that $F_{\mathrm{s}}\ll F_{\mathrm{d}}$. This is attributed to the large difference between the non-equilibrium molecular occupancy and that of the substrate sites around the interaction region, as we discuss in the next section [see Fig. 2(a)]. Therefore, throughout the rest of this work, we will restrict our analysis to the intramolecular term $F_{\mathrm{d}}$, defined through the following force operator:
@@
[…]
$\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)=\bm{H}_{\mathrm{dd}}+\bm{H}_{\mathrm{ds}}\left(\epsilon\bm{I}-\bm{H}_{\mathrm{ss}}\right)^{-1}\bm{H}_{\mathrm{sd}},$ (17)
where the diagonal blocks $\bm{H}_{\mathrm{dd}}$ and $\bm{H}_{\mathrm{ss}}$ are the bare molecular and substrate Hamiltonian matrices of Eqs. (3) and (9), respectively, while the off-diagonal blocks $\bm{H}_{\mathrm{ds}}$ and $\bm{H}_{\mathrm{sd}}$ are the coupling matrices between these subsystems, representing the interaction Hamiltonian of Eq. (7). By taking $\epsilon$ as a fixed parameter, we diagonalize $\tilde{\bm{H}}_{\mathrm{dd}}$ according to:
$\tilde{\bm{H}}_{\mathrm{dd}}(\epsilon)\bm{u}_{m}(\epsilon)=\lambda_{m}(\epsilon)\bm{u}_{m}(\epsilon),$ (18)
and obtain the substrate-renormalized molecular eigenstates $\bm{u}_{m}$. Notice that we are not solving here the (non-linear) stationary Schrödinger equation which leads to the poles of the Green’s function through the solutions of $\lambda_{m}(z)=z$, with $z\in\mathbb{C}$. Instead, we use the transformation matrix $\bm{U}$ that diagonalizes $\tilde{\bm{H}}_{\mathrm{dd}}$ to compute the molecule’s density matrix through the kernel:
@@
where $\bm{\Pi}_{\mathrm{d}}=\bm{\Pi}^{\dagger}_{\mathrm{d}}$ is the projection operator matrix in the molecular subspace. We will call these molecular states ‘bonding’ (b) and ‘antibonding’ (a), since they essentially preserve the properties of the symmetric and antisymmetric states defined in Eq. (6), for the considered regimes of parameters.
@@ III.1 Non-equilibrium molecular destabilization
We initiate our analysis by considering a pristine armchair-edged graphene nanoribbon defined by $N$ carbon atoms along the transversal direction [40] [see Fig. 1(b)]. In particular, we set $N$ under the condition $N=3\ell+2$, with $\ell\in\mathbb{N}$, in order to obtain a metallic ribbon. For this analysis we use $\ell=25$, resulting in a nanoribbon width of approximately $W=\sqrt{3}(N-1)a/2\approx 9.3$ nm. The molecular center is located at a distance of $1.5$ Å above the substrate plane, with its bond axis aligned along $\theta=\pi/2$, i.e., parallel to the $y$ direction. In Appendix B we extended the analysis to various molecular orientations, though no qualitative differences appear for both the molecular occupancies and the intramolecular force. Additionally, its center is located at the lower edge, i.e., $y_{0}=(1-N)\sqrt{3}a/4$, measured from the center of the ribbon. The two atoms are separated by a distance of $1.5$ Å from each other along the bond axis, such that the radius of the molecule is $R=0.75$ Å. For the topological case ($\gamma_{1}\neq 0$), this setup allows us to probe the interaction between the molecule and the localized edge-states. To ensure that the edge states are well defined and spatially separated, the ribbon width must be chosen to be larger than the localization length to avoid hybridization between opposite edges.
Figure 2: Molecular destabilization under non-equilibrium conditions. (a) Density of states (DOS), using the logarithmic scale $\log(1+\mathcal{N}_{q})$, projected on the substrate region around the molecule ($q=\mathrm{s}$, blue) and on the molecule itself ($q=\mathrm{d}$, red) for graphene (G, left panels) and Kane-Mele (KM, right panels) substrates. (b) Occupation probabilities of the bonding (solid lines) and antibonding (dotted lines) levels as functions of the bias voltage for graphene (G, blue lines) and Kane-Mele model (KM, red lines) substrates. (c) Corresponding non-equilibrium intramolecular force, according to Eq. (16). We use a relaxation energy $\eta=10^{-5}\gamma_{0}$, an equilibrium chemical potential $\mu_{0}=0$, and $k_{\mathrm{B}}T=10^{-3}\,\gamma_{0}$.
@@
With these parameters, in our system the bonding level lies below the Dirac point ($\epsilon=0$), and the antibonding level above it. If the molecule were isolated, both levels would be symmetrically placed with respect to the Dirac point. However, due to hybridization with the substrate, this balance is broken and the bonding level shifts more than the antibonding one, ending up closer to the Dirac point. This can be seen in Fig. 2(a), where we show the projected density of states (DOS), calculated as:
$\mathcal{N}_{q}(\epsilon)=\sum_{\alpha}\sum_{i\in q}[\bm{\mathcal{K}}^{\rho}_{\alpha}]_{ii}(\epsilon),$ (21)
where $q=\{\mathrm{s},\mathrm{d}\}$ denotes the projection region. We use the logarithmic scale $\log(1+\mathcal{N}_{q})$ to reduce the high contrast produced by the bonding and antibonding peaks. The s region includes the 36 closest sites surrounding the molecule, while the d region consists only of the two molecular sites. For the d-projected DOS (on the molecule), we observe the shifted bonding peak below $\epsilon=0$ and the antibonding peak above it. Due to the molecular hybridization with the substrate, these peaks also appear in the s-projected DOS, alongside the typical van Hove singularities of the ribbon. Regarding the magnitude, the d-projected DOS peaks are about 135 times larger than the s-projected DOS for both graphene and Kane-Mele substrates.
The large difference between the molecular and substrate densities determines a perturbative regime for the effective interaction between these subsystems. Under non-equilibrium conditions, this implies a small $F_{\mathrm{s}}$ compared with the intramolecular force.
In addition, within the framework of Refs. [41, 42, 43], we estimate that Landauer resistivity dipoles [44], related with the current-induced charge redistribution around the molecule, can be disregarded in a first approximation due to the observed unbalance in the projected DOS.
@@
In Fig. 2(b) we show the occupation probability of the molecular levels as a function of the applied bias for both trivial graphene ($\gamma_{1}=0$) and topological ($\gamma_{1}\neq 0$) substrates. At equilibrium ($V=0$), the bonding level shows its maximum occupancy while the antibonding level is unoccupied. This is because we set the equilibrium chemical potential at the Dirac point, i.e., $\mu_{0}=0$, and the bonding (antibonding) level lies below (above) this energy point. When a bias voltage is applied and increased, the occupancy of the bonding level starts to decrease while the antibonding level becomes progressively populated. In this case, the molecular levels enter the region known as the “bias window”—that is, between the two chemical potentials. Here, a finite electron current flows through the molecule, thereby modifying the previous equilibrium occupancies.
@@
As can be seen, the variations in the bonding and antibonding occupancies take place at different voltage values: the bonding level depopulates first, while the antibonding occupancy begins to increase at slightly higher voltages. This difference arises from the asymmetric energy alignment of the bonding and antibonding levels with respect to $\mu_{0}$, as previously discussed. Importantly, both the depopulation of the bonding level and the subsequent population of the antibonding level contribute to the destabilization of the molecule. In particular, for large biases the antibonding occupancy can even surpass that of the bonding level. These results indicate that the interaction with either graphene or Kane-Mele substrates can drive molecular dissociation through electron transport.
To account for additional relaxation processes due to the environment, we introduce a small imaginary part $\eta$ in the energy variable, i.e., $\epsilon\rightarrow\epsilon+i\eta$. This would correspond, e.g., to coupling the system to a phonon bath that allows electrons to decay into lower energy levels. The value of $\eta$ effectively sets a timescale for the electronic relaxation: a smaller $\eta$ corresponds to a longer lifetime of electronic states, while a larger $\eta$ implies faster relaxation. This allows us to capture the effect of finite-lifetime processes, such as electron-phonon interactions, without the need to model them explicitly (see Appendix A). As a consequence of this relaxation, the molecular occupancies are not necessarily normalized. These deviate from their ideal values because we impose a finite relaxation time through the $\eta$ parameter which, in turn, reflects the decay rates to the contacts. For example, the kernel $\bm{\mathcal{K}}_{\mathrm{L}}^{\rho}]_{\mathrm{b},\mathrm{b}}$, related to the bonding level’s occupation from the left contact, develops a Lorentzian peak shape due to the molecule-substrate hybridization. Its width is characterized by the sum of the rate $\Gamma_{\mathrm{L},\mathrm{b}}$, associated with the molecular hybridization with the left contact, and $\eta$. Energy integration of this kernel [cf., Eq. (13)] yields a factor $\Gamma_{\mathrm{L},\mathrm{b}}/(\Gamma_{\mathrm{L},\mathrm{b}}+\eta)$ which bounds the level occupancy below its ideal value. Therefore, the $\eta$ parameter establishes an additional decay mechanism that competes with that related with the coupling to the contacts. Indeed, if we set $\eta=0$, the level occupancy recovers its ideal value, regardless of the value of $\Gamma_{\mathrm{L},\mathrm{b}}$.
Indeed, this relaxation model could be improved. For example, in our tight-binding description of the composite system, we could implement a decoherence time or length using the D’Amato and Pastawski model [45]. We estimate that such an inelastic scattering model would also yield a characteristic time scale to be compared with the molecular occupation time from the contacts.
@@
It is often more intuitive to analyze the destabilization of the dimer in terms of the non-equilibrium force acting on the molecular bond. For this purpose we show, in Fig. 2(c), the non-equilibrium component of the intramolecular force of Eq. (16). In the absence of an applied bias, the non-equilibrium force is zero, which signals molecular stability. As the bias increases, the force exhibits a sharp rise, followed by a small shoulder over a narrow range of $V$ before rising again and reaching its maximum value at higher biases. Once more, this non-monotonic behavior, particularly the presence of the intermediate plateau, can be attributed to the shifts of the molecular levels when coupled with the substrate. Consequently, the initial increase in force is associated with the depletion of the bonding occupancy, while the subsequent rise is driven by the occupation of the antibonding level.
@@
It may be useful to establish a comparison between the equilibrium and non-equilibrium components of the force when the two molecular levels lie within the bias window. By relating the intramolecular force with the molecular occupancies through (see Appendix A): [^1] Notice that for the considered parameters, the bonding and antibonding occupancies can be used here, provided the small mixing between the symmetric and antisymmetric states due to the substrate.
$F_{\mathrm{d}}\approx 2\frac{\mathrm{d}\gamma_{\mathrm{d}}}{\mathrm{d}R}(\rho_{\mathrm{b}}-\rho_{\mathrm{a}}),$ (22)
@@
we obtain, for $eV>2\gamma_{\mathrm{d}}$, that $|F^{\mathrm{ne}}/F^{\mathrm{eq}}|\approx(1+\rho_{\mathrm{a}}/\rho_{\mathrm{b}})/2$. For the considered substrates we find that $\rho_{\mathrm{a}}>\rho_{\mathrm{b}}$, which implies that the destabilizing non-equilibrium component is larger than the stabilizing equilibrium term. In addition, it can be seen that $F_{\mathrm{d},\mathrm{KM}}^{\mathrm{ne}}>F_{\mathrm{d},\mathrm{G}}^{\mathrm{ne}}$, indicating a higher dissociation efficiency in the topological substrate.
@@
In this section, we analyze the effect of the ribbon width on the molecular occupation, considering the extended nature of metallic states in graphene-based substrates and the localized nature of edge states in topological substrates. This analysis is particularly relevant in systems with coupling to an external environment—modeled via a finite broadening parameter $\eta$—that introduces escape channels that affect the molecular occupancies. A very small $\eta$ would act as a mathematical regularization which, for energy levels below the Fermi energy, yields occupancies close to its ideal value regardless of the system size. A finite $\eta$, on the other hand, allows the ribbon’s spatial characteristics to play a physical role in the transport processes, since the decay rates $\Gamma_{\alpha,\mathrm{b/a}}$ modulate the molecular occupancies. This effect is illustrated in Appendix C.